5.2 Paris to Moscow. On your summer study abroad program in Europe you stay an extra...

5.2

Paris to Moscow. On your summer study abroad program in Europe you stay an extra two weeks to travel from Paris to Moscow. You leave Paris with 2 comma 200 euros (euro or EUR) in your belt pack. Wanting to exchange all of these for Russian rubles (R or RUB), you obtain the following quotes:

Spot rate

USDequals=1.00EUR (or $/euro€ )                                         1.1187

Spot rate RUBequals=1.00 USD (or R/$)                                       61.96

a. What is the Russian ruble to euro cross rate?

b. How many Russian rubles will you obtain for your euros?
Answer #1

a

Rouble euro cross = USD/EUR*RUB/USD = 1.1187*61.96=69.314

b

Amount in RUB = EURO amount *cross rate = 2200*69.314=152490.8